Providing for Doubtful Accounts

At the end of the current year, the accounts receivable account has a debit balance of $1,154,000 and sales for the year total $13,090,000.

  1. The allowance account before adjustment has a credit balance of $15,600. Bad debt expense is estimated at 3/4 of 1% of sales.
  2. The allowance account before adjustment has a credit balance of $15,600. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $49,900.
  3. The allowance account before adjustment has a debit balance of $6,400. Bad debt expense is estimated at 1/4 of 1% of sales.
  4. The allowance account before adjustment has a debit balance of $6,400. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $53,100.

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. $
b. $
c. $
d. $
Answer #1

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. 13090000*1%*3/4 = 98175
b. $49900-15600 = 34300
c. $13090000*1$*1/4 = 32725
d. $53100+6400 = 59500